a line make intercept a and b on the coordinate axes.now axis is rotated by angle x.now that line make equal intercept on the new coordinate axis.what is the value of tanx

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Manoj

equation of line in first co ordinate system

x/a  +y/b =i

let new coordinaate system are x' and y'

so x' =xcosΘ +ysinΘ

   y' = ycosΘ -xsinΘ

so equation of line in new co ordinate system

x'/c +y'c =1

now put value of x' and y'

x(cosΘ -sinΘ)/c  + y( sinΘ + cosΘ)/c =1

these two line are same

so  (cosΘ -sinΘ)  =c/a

    ( sinΘ + cosΘ) =c/b

 solve these two

tan Θ =(a-b)/(a+b)

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin